N(x) = -0.4x^2 +8x + 12 IS the formula. What day will the ticket sales peak on? and how many tickets will be sold on that day.
Answer by stanbon(75887)

N(x) = -0.4x^2 +8x + 12 IS the formula. What day will the ticket sales peak on? and how many tickets will be sold on that day.
------
a = -0.4 ; b = 8 ; c = 12
---
Max tickets sales occurs when x = -b/(2a) = -8/(2*-0.4) = 10
---
Max ticket sales = N(10) = -0.4^10^2 + 8*10 +12
= -40 + 80 + 12
= 52
==========
Cheers,
Stan H.